# Graph-Structured Topic Modeling for Documents with Spatial or Covariate Dependencies

### v1

@@ 2.3 Estimation procedure: GpLSI
Theoretical insights from Klopp et al. (2021) help explain why topic modeling deteriorates in low-$N$ regimes. When the number of words per document is too small, the observed frequency matrix $X$ can be viewed as a highly noisy approximation of $M$, causing the estimated singular vectors $\hat{U}$ to deviate significantly from the true underlying point cloud $U$. To mitigate this issue, Klopp et al. (2021) suggest a preconditioning step that improves the estimation of the singular vectors in noisy settings.
In this paper, we take a different approach by exploiting the graph structure associated with the documents to reduce the noise in $X$. Rather than preconditioning the empirical frequency matrix, we propose an additional denoising step that leverages the graph structure to produce more accurate estimates of $U$, $V$, and $\Lambda$.
We aim to accurately estimate the topic mixture matrix $W$ under Assumption 1 by adapting the procedure of Klopp et al. (2021). Specifically, we consider a four-step algorithm for estimating both $W$ and $A$. Compared to the original approach described in Section 2.2, our adaptation modifies Step 1 and Step 4:
@@
We replace the first step of Section 2.2 with a graph-aligned SVD of the empirical word-frequency matrix $X$. More specifically, in the graph-aligned setting, we assume that the underlying frequency matrix $M$ belongs to the set:
@@
2.3.0.1 Iterative Graph-Aligned SVD
We propose a method based on power-iterations for Step 1 that iteratively updates the left and right singular vectors while constraining the left singular vector to be aligned with the graph (Algorithm 1). A similar approach has already been studied under Gaussian noise in Yang et al. (2016) where sparsity constraints were imposed on the left and right singular vectors.
Drawing inspiration from this method and adapting it to the multinomial noise setting, Algorithm 1 iterates between three steps. The first consists of denoising the left singular subspace by leveraging the graph-smoothness assumption (Assumption 1). At iteration $t$, we define:
@@ 2.3.0.1 Iterative Graph-Aligned SVD
$\mathrm{SVD}(\bar{U}^{t})=\widehat{U}^{t}$
As shown in the next subsection, this algorithm can be proven to considerably improve upon the original SVD estimates. Denoting the final estimates after $t_{\max}$ iterations as $\widehat{U},\widehat{V}$, these estimates can then be plugged into Steps 2-4 to estimate $W$ and $A$. The improved estimation of $\widehat{U},\widehat{V}$ can then be shown to translate subsequently into a more accurate estimation of the matrices $W$ and $A$ (Theorems 3 and 4 presented in the next section).
Although our theoretical results depend on choosing an appropriate level of regularization $\hat{\rho}^{t}$, the theoretical value of $\hat{\rho}^{t}$ depends on unknown graph quantities. In practice, therefore, we need a cross validation technique to choose the optimal $\hat{\rho}^{t}$ in each iteration. We devise a novel graph cross-validation method which effectively finds the optimal graph regularization parameter by partitioning nodes into folds based on a natural hierarchy derived from a minimum spanning tree. The procedure is summarized in Algorithm 2.
Remark 3
@@
5:     2. SVD of $\bar{U}^{t}$:
$\widehat{U}^{t}\leftarrow\text{Left Singular Vectors in }SVD_{K}(\bar{U}^{t})$
6:     3. SVD of $\widehat{V}^{t}$:
$\widehat{V}^{t}\leftarrow\text{Left Singular Vectors in }SVD_{K}(X^{\top}\widehat{U}^{t})$
7:     4. Calculate score $s=\lVert\widehat{P}_{u}^{t}X\widehat{P}_{v}^{t}-\widehat{P}_{u}^{t-1}X\widehat{P}_{v}^{t-1}\rVert$, $\widehat{P}_{u}=\widehat{U}^{t}(\widehat{U}^{t})^{\top}$, $\widehat{P}_{v}=\widehat{V}^{t}(\widehat{V}^{t})^{\top}$
@@
Remark 4
The penalty in Equation (10) is known as the total-variation penalty in the computer vision literature. As noted in Hütter and Rigollet (2016), this type of penalty is usually a good idea whenever the rows of $W$ take similar values, or may at least be well approximated by piecewise-constant functions. In the implementation of our algorithm, we employ the solver of Sun et al. (2021), as it is the most efficient algorithm available for this type of problem.
2.3.0.2 Optimization of ${\hat{\rho}^{t}}$
We propose using a cross validation technique that leverages a minimum spanning tree (MST-CV) to find the optimal $\hat{\rho}^{t}$ in Equation (10). Conventional cross validation techniques sample either nodes or edges to divide the dataset into folds. However, these approaches can disrupt the graph structure and underestimate the strength of the connectivity of the graph. We instead devise a new rule for dividing documents into folds using a minimum spanning tree. This technique is an extension of the cross-validation procedure proposed by Tibshirani and Taylor (2012) for the line graph. Given a minimum spanning tree $\mathcal{T}$ of $\mathcal{G}$, we randomly choose a source document $d_{s}$. For each document $d_{i}$, we calculate the shortest path distance $\mathrm{d}_{\mathcal{T}}(d_{i},d_{s})$. Note that this distance is always an integer. We divide the documents into $b$ folds based on the modulus of their distance from the source node: $\mathrm{d}_{T}(d_{i},d_{s})\mod b$. Through this construction of folds, we can ensure that for every document, at least one of its 1-hop neighbors is in a different fold.
Let $X_{i\cdot}$ be the $i^{th}$ row of $X$. For each leave-out fold $\mathcal{I}_{k}$, $k\in[b]$, we interpolate the corresponding documents $X_{i\cdot}$$\forall i\in\mathcal{I}_{k}$, filling the missing document information with the average of corresponding neighbors in $\mathcal{I}_{k}^{C}$. This prevents us from using any information from the leave-out fold in training when calculating the cross-validation error. For each $\rho$ in the given grid $\rho_{[1:r]}=\{\rho_{1},\rho_{2},\cdots,\rho_{r}\}$, we use the interpolated observations $X^{k}$ to solve the optimization problem in Equation (10) and estimate $\widehat{X}^{k}_{\mathcal{I}_{k}}$, the submatrix of $\widehat{X}^{k}$ with rows in index set $\mathcal{I}_{k}$, to calculate cross-validation error. The method is described in Algorithm 2.
2.3.0.3 Initialization
@@
for some constants $C$ and $C^{*}>0$. The proof of the theorem is provided in Appendix A.

### v2

@@ 2.3 Estimation procedure: GpLSI
Theoretical insights from Klopp et al. (2021) help explain why topic modeling deteriorates in low-$N$ regimes. When the number of words per document is too small, the observed frequency matrix $X$ can be viewed as a highly noisy approximation of $M$, causing the estimated singular vectors $\widehat{U}$ to deviate significantly from the true underlying point cloud $U$. To mitigate this issue, Klopp et al. (2021) suggest a preconditioning step that improves the estimation of the singular vectors in noisy settings.
In this paper, we take a different approach by exploiting the graph structure associated with the documents to reduce the noise in $X$. Rather than preconditioning the empirical frequency matrix, we propose an additional denoising step that leverages the graph structure to produce more accurate estimates of $U$, $V$, and $\Lambda$. Specifically, we modify the SVD of $X$ in Step 1 and estimation of topic matrix $A$ in Step 4 described in Section 2.2.
@@
We replace Step 1 of Section 2.2 with a graph-aligned SVD of the empirical word-frequency matrix $X$. More specifically, in the graph-aligned setting, we assume that the underlying frequency matrix $M$ belongs to the set:
@@
Iterative Graph-Aligned SVD.
We propose a power iteration method for Step 1 that iteratively updates the left and right singular vectors while constraining the left singular vector to be aligned with the graph (Algorithm 1). A similar approach has already been studied under Gaussian noise in Yang et al. (2016) where sparsity constraints were imposed on the left and right singular vectors.
Drawing inspiration from Yang et al. (2016) and adapting this method to the multinomial noise setting, Algorithm 1 iterates between three steps. The first consists of denoising the left singular subspace by leveraging the graph-smoothness assumption (Assumption 1). At iteration $t$, we solve:
@@
5:     2. SVD of $\bar{U}^{t}$: $\widehat{U}^{t}\leftarrow\text{Left Singular Vectors in }SVD_{K}(\bar{U}^{t})$
6:     3. SVD of $\widehat{V}^{t}$: $\widehat{V}^{t}\leftarrow\text{Left Singular Vectors in }SVD_{K}(X^{\top}\widehat{U}^{t})$
7:     4. Calculate the score $s=\lVert\widehat{P}_{u}^{t}X\widehat{P}_{v}^{t}-\widehat{P}_{u}^{t-1}X\widehat{P}_{v}^{t-1}\rVert$, $\widehat{P}_{u}=\widehat{U}^{t}(\widehat{U}^{t})^{\top}$, $\widehat{P}_{v}=\widehat{V}^{t}(\widehat{V}^{t})^{\top}$
@@
Denoting the final estimates after $t_{\max}$ iterations as $\widehat{U},\widehat{V}$, these estimates can then be plugged into Steps 2-4 to estimate $W$ and $A$. The improved estimation of $\widehat{U},\widehat{V}$ can be shown to translate into a more accurate estimation of the matrices $W$ and $A$ (Theorems 3 and 4 presented in the next section).
Although our theoretical results depend on choosing an appropriate level of regularization $\hat{\rho}^{t}$, the theoretical value of $\hat{\rho}^{t}$ depends on unknown graph quantities. In practice, therefore, the optimal $\hat{\rho}^{t}$ must be chosen in each iteration using cross-validation. We devise a novel graph cross-validation method which effectively finds the optimal graph regularization parameter by partitioning nodes into folds using a minimum spanning tree. We defer the detailed procedure to Section A of the Appendix.
Remark 3.
Remark 4.
The penalty $\|\Gamma U\|_{21}$ is known as the total-variation penalty in the computer vision literature. As noted in Hütter & Rigollet (2016), this type of penalty is usually a good idea whenever the rows of $W$ take similar values, or may at least be well approximated by piecewise-constant functions. In the implementation of our algorithm, we employ the solver of developed by Sun et al. (2021), as it is the most efficient algorithm available for this type of problem.
Initialization.
@@
for some constants $C$ and $C^{*}>0$.
The proof of the theorem is provided in Section B of the Appendix.
3 Theoretical results
3.1 Denoising the singular vectors
The improvement in the estimation of the singular vectors induced by our iterative denoising procedure is characterized in the following theorem.
Theorem 2.
Let Assumption 1 to 5 hold. Assume $\max(K,p)\leq n$ and $\sqrt{K}\leq p$. Denote $\widehat{U},\widehat{V}$ as outputs of Algorithm 1 after $t_{\max}$ iterations. If $N$ satisfies
The proof of this result is provided in Section C of the Appendix.
Remark 5.
This result is to be compared against the rates of the estimates obtained without any regularization. In this case, the results of Klopp et al. (2021) show that with probability at least $1-o((n+p)^{-1})$, the error in (13) is of the order of $O(K\sqrt{\log(n)/{N}})$. Both rates thus share a factor $K\sqrt{\log(n)/N}$. However, the spatial regularization in our setting allows us to introduce an additional factor of the order of $\frac{1}{\sqrt{n}}(\sqrt{n_{\mathcal{C}}}+\rho(\Gamma)\sqrt{s\lambda_{\max}(\Gamma)})$. The numerator in this expression can be interpreted as the effective degrees of freedom of the graph, and for disconnected graphs ($\lambda_{\max}(\Gamma)=0,n_{\mathcal{C}}=n$), the results are identical. However, for other graph topologies (e.g. the 2D grid, for which $\lambda_{\max}(\Gamma)$ is bounded by a constant, $\rho\lesssim\log(n)$ (see Hütter & Rigollet (2016)) and $n_{\mathcal{C}}=1$), our estimator can considerably improve the estimation of the singular vectors (see Section 3.3).
3.2 Estimation of $W$ and $A$
We now show how our denoised singular vectors can be used to improve the estimation of the mixture matrix $W.$
Theorem 3.
Let Assumptions 1 to 5 hold. Let $\rho(\Gamma),s,n_{\mathcal{C}}$, and $n_{\mathcal{C}_{\min}}$ be given as (3)-(5). Assume $\max(K,p)\leq n$ and $\sqrt{K}\leq p$. Let $\widehat{W}$ denote the estimator of the mixture matrix (Equation 2) obtained by running the SPA algorithm on the denoised estimates of the singular vectors (Algorithm 1). If $N$ satisfies the condition stated in (12), then there exists a constant $C>0$ such that with probability at least $1-o(n^{-1})$,
where $\mathcal{P}$ denotes the set of all permutations.
Theorem 3 shows that $\widehat{W}$ is highly accurate as long as the document lengths are large enough, as defined by $N\gtrsim({n_{\mathcal{C}}}+\rho^{2}(\Gamma){s\lambda_{\max}(\Gamma)})\log(n)/n$. This requirement is more relaxed than the condition $N\gtrsim\log(n+p)$ for pLSI provisioned in Theorem 1 and Corollary 1 of Klopp et al. (2021). This indicates that GpLSI can produce accurate estimates even for smaller $N$, by sharing information among neighboring documents. The shrinkage of error due to graph-alignment is characterized by the term $\frac{1}{\sqrt{n}}(\sqrt{n_{\mathcal{C}}}+\rho(\Gamma)\sqrt{s\lambda_{\max}(\Gamma)})$, which is equal to one when the graph $\mathcal{G}$ is empty. In general, the effect of the regularization depends on the graph topology. Hütter & Rigollet (2016) show in fact that the inverse scaling factor verifies: $\rho(\Gamma)\leq\sqrt{2}/\sqrt{\lambda_{n-1}(L)}$. The quantity $\lambda_{n-1}(L)$, also known as the algebraic connectivity, provides insights into the properties of the graph, such as its connectivity. Intuitively, higher values of $\lambda_{n-1}(L)$ reflect more tightly connected graphs (Chung 1997), thereby reducing the effective degrees of freedom induced by the graph-total variation penalty. By contrast, $\lambda_{\max}(\Gamma)$ can be coarsely bounded using the maximum degree $d_{\max}$ of the graph: $\lambda_{\max}(\Gamma)\leq\sqrt{2d_{\max}}$ (Anderson Jr & Morley 1985, Zhang 2011). Consequently, we can expect our procedure to work well on well-connected graphs with bounded degree. Examples include for instance grid-graphs, $k$-nearest neighbor graphs, or spatial (or planar) graphs. We provide a more detailed discussion and more explicit bounds for specific graph topologies in Section 3.3.
Furthermore, using the inequality $\|\widehat{W}-WP\|_{11}\leq\sqrt{Kn}\|\widehat{W}-WP\|_{F}$, it immediately follows that:
Corollary 1.
$\min_{P\in\mathcal{P}}\|\widehat{W}-WP\|_{11}\leq CK^{3/2}\sqrt{\frac{n\log(n)}{N}}\left(\sqrt{n_{\mathcal{C}}}+\rho(\Gamma)\sqrt{s\lambda_{\max}(\Gamma)}\right).$ (15)
where $\mathcal{P}$ denotes the set of all permutations.
Finally, we characterize the error bound of $\widehat{A}$. The full proofs of Theorems 3 and 4 are deferred to Section D of the Appendix.
Theorem 4.
where, denoting $P$ the permutation matrix that minimises the distance between $\widehat{W}$ and $W$ in (14), we take $\tilde{P}$ to be its inverse: $\tilde{P}=P^{-1}$.
Remark 6.
The previous error bound of $A$ indicates that the accuracy of $\widehat{A}$ primarily relies on the accuracy of $\widehat{W}$, which is to be expected, since $A$ is estimated by regressing $X$ on the estimator $\widehat{W}$. While the error rate may not achieve the minimax-optimal rate $C\sqrt{p/(nN)}$ derived in Ke & Wang (2017), we found that this procedure is more accurate than the estimator $\widehat{A}$ proposed in Klopp et al. (2021), as confirmed by synthetic experiments in Section 3.4.
3.3 Refinements for special graph structures
We now analyze the behavior of the error bound provided in Theorem 3 for different graph structures, further expliciting the dependency of our bounds on graph properties.
Erdös-Rényi random graphs.
We first assume that the graph $\mathcal{G}$ is an Erdös-Rényi random graph where each pair of nodes is connected with probability $p=p_{0}\frac{\log(n)}{n}$ for a constant $p_{0}>1$. In this case, Hütter & Rigollet (2016) show that with high probability, the algebraic connectivity $\rho(\Gamma)$ is of order $O(\frac{1}{\log(n)})$. Moreover, the maximal degree is of order $\log(n)$ and the graph is almost surely connected (Van Der Hofstad 2024). Under this setting, the error associated to our estimator $\widehat{W}$ becomes:
$\min_{P\in\mathcal{P}}\|\widehat{W}-WP\|_{F}\leq C_{1}K\sqrt{\frac{\log(n)}{N}}\left(1+s^{\frac{1}{2}}\log(n)^{-\frac{3}{4}}\right).$ (17)
Grid graphs.
We also derive error bounds for grid graphs, which are commonly occurrences in the analysis of spatial data and for applications in image processing:
2D grid graph:
Let $\mathcal{G}$ be a 2D grid graph on $n$ vertices. Hütter & Rigollet (2016) show that, in that case, the inverse scaling factor is such that $\rho(\Gamma)\lesssim\sqrt{\log(n)}$ The error of our estimator thus becomes:
$K$-grid graph, $k\geq 3$:
In this case, Hütter & Rigollet (2016) show that the inverse scaling factor is bounded by a constant $C(k)$, that depends on the dimension $k$ but is independent of $n$. In this case, the error of our estimator becomes:
$\min_{P\in\mathcal{P}}\|\widehat{W}-WP\|_{F}\leq CK\sqrt{\frac{\log(n)}{N}}\left(1+\sqrt{s}\right)\leq C_{3}K\sqrt{\frac{s\log(n)}{N}}.$ (19)
3.4 Synthetic Experiments
Experimental Protocol
To generate documents, we sample $n$ points uniformly on unit square $[0,1]^{2}$, and cluster them into ${n_{grp}}=30$ groups using a simple k-means algorithm. For each group, we generate the topic mixture as $\mathbf{\alpha}\sim\mathrm{Dirichlet}(\mathbf{u})$ where $u_{k}\sim\mathrm{Unif}(0.1,0.5)$ ($k\in[K]$). Small random noise $N(0,0.03)$ is added to $\mathbf{\alpha}$ for each document in the group, and we permute it so that the biggest element of $\alpha$ is assigned to the group’s predominant topic. $A$ is generated by sampling each entry $A_{kj}$ from a uniform distribution, normalizing each row to ensure that $A$ is a stochastic matrix. A detailed description of the data generating process is provided in Section F of the Appendix.
To assess the performance of GpLSI, we compare it against several established methods, including the original pLSI algorithm proposed by Klopp et al. (2021), TopicSCORE (Ke & Wang 2017), LDA (Blei et al. 2001), and the Spatial LDA (SLDA) approach of Chen et al. (2020). In addition, to highlight the efficiency of our iterative algorithm, we present results from a baseline variant that employs only a single denoising step. This one-step method is described in greater detail in Section C of the Appendix. To implement these algorithms, we use the R package TopicScore and the LDA implementation of the Python library sklearn. For SLDA, we use of the Python package spatial-lda with the default settings of the algorithm. We run 50 simulations and record the $\ell_{1}$ error, $\ell_{2}$ error of $W$ and $A$, and the computation time across various parameter settings $(p,N,n,K)$, reporting medians and interquartile ranges. To evaluate the performance of methods in difficult scenarios where document length $N$ is small compared to vocabulary size $p$, we check the errors for different combinations of $N=10,30,50,100,200,1000$ and $p=20,30,50,100,200,500$.
Results
Figure 1 demonstrates that GpLSI achieves the lowest $\ell_{2}$ error for $W$, even in scenarios with very small $N$. This shows that sharing information across similar documents on a graph improves the estimation of topic mixture matrix. Notably, while LDA and pLSI exhibit modest performance, they fail in regimes where $N<100$ and $p\geq 200$. We also confirm that the one-step denoising variant of our method achieves a lower error estimation error than pLSI and LDA in settings where $N<<p$.
Figure 2: $\ell_{2}$ error of $W$ (left) and $A$ (middle) and computation time (right) for different corpus size $n$ and number of topics $K$. Here, $N=30$ and $p=30$. Errors are normalzied by $n$.
We also examine how the estimation errors scale with the corpus size $n$ and the number of topics $K$, as shown in Figure 2. We observe that GpLSI substantially outperforms other methods, particularly for the estimation of $W$. GpLSI achieves the lowest error for $A$, as we show in Section F of the Appendix. Similar patterns hold for $\ell_{1}$ errors of $A$ and $W$ also provided in Section F of the Appendix.
4 Real-World Experiments
4.1 Tumor Microenvironment discovery: the Stanford Colorectal Cancer dataset
We first consider the analysis of CODEX data, which allows the identification of individual cells in tissue samples, holding valuable insights into cell interaction profiles, particularly in cancer, where these interactions are crucial for immune response. Since cellular interactions are hypothesized to be local, these patterns are often referred to as “tumor microenvironments”. In the context of topic modeling, we can regard a tumor microenvironment as a document, immune cell types as words, and latent characteristics of a microenvironment as a topic. However, due to the small number of words per document, the recovery of the topic mixture matrix and the topics themselves can prove challenging. Chen et al. (2020) propose using the adjacency of documents to assign similar topic proportions to neighboring tumor cells. Similarly, we construct a spatial graph based on proximity of tumor microenvironments to uncover novel tumor-immune cell interaction patterns.
The first CODEX dataset is a collection of 292 tissue samples from 161 colorectal cancer patients collected at Stanford University (Wu et al. 2022). The locations of the cells were retrieved using a Voronoi partitioning of the sample, and the corresponding spatial graphs were constructed encoding the distance between microenvironments. More specifically, we define a tumor microenvironment as the 3-hop neighborhood of each cell, following the definition originally used by Wu et al. (2022). Each microenvironment contains 10 to 30 immune cells of 8 possible types. This aligns with the setting where the document length $N<30$ is small compared to the vocabulary size $p=8$.
Figure 3: (A) Estimated tumor-immune topic weights of GpLSI, pLSI, and LDA. Topic weights are aligned across methods using cosine similarity. (B) Topic alignment paths of GpLSI, pLSI, and LDA using R package alto. (C) Pairwise $\ell_{1}$ distance and cosine similarity of topic weights from different batches of patients.
We aggregate frequency matrices and tumor cellular graphs of all samples and fit three methods: our proposed GpLSI, the original pLSI approach of Klopp et al. (2021), and LDA (Blei et al. 2001) to estimate tumor-immune topics. The estimated topic weights of $K=6$ are illustrated in Figure 3(A). After aligning topic weights across methods, we observe similar immune topics (Topic 1, 2, 3) in GpLSI and LDA which are not found in the estimated topics of pLSI.
To determine the optimal number of topics $K$, we use the method proposed by Fukuyama et al. (2023). In this work, the authors construct “topic paths” to track how individual topics evolve, split or merge, as the number of topics $K$, increases. We observe in Figure 3(B) that while the GpLSI path has non-overlapping topics up to $K=6$, other methods fail to provide consistent and well-separated topics.
To evaluate the quality and stability of the recovered topics, we also measure the coherence of the estimated topic weights of batches of samples, as suggested in Tran et al. (2023). We divide 292 samples into five batches and estimate the topic weights $A^{b}$ for $b\in[5]$. For every pair of batches $(b,b^{{}^{\prime}})$, we align $A^{b}$ and $A^{b^{{}^{\prime}}}$ (we permute $A^{b^{{}^{\prime}}}$ with $P$ where $P=\arg\min_{P\in\mathcal{P}}\|A^{b}-PA^{b^{{}^{\prime}}}\|$) and measure the entry-wise $\ell_{1}$ distance and cosine similarity. We repeat this procedure five times and plot the scores in Figure 3(C). We notice that GpLSI provides the most coherent topics across batches for $K=3,4,5$. Combining with the metrics of LDA, we can choose the optimal $K$ as 5 or 6.
Figure 4: (A) AUC for predicting cancer recurrence using isometric log-ratio transformed topic proportions (top) and dichotomized topic proportions (bottom) as covariates. (B) Kaplan-Meier curves based on dichotomized topic proportions using GpLSI.
Next, we conduct survival analysis to identify the immune topics associated with higher risk of cancer recurrence. We consider two logistic models with different covariates to predict cancer recurrence and calculate the area under the curve (AUC) of the receiver-operating characteristic (ROC) curves to evaluate model performance.
In the first model, we use the proportion of each microenvironment topic as covariates for each sample. Since the $K$ covariates sum up to one, we apply isometric log-ratio transformation to represent it with $K-1$ orthonormal basis vectors In the second model, we dichotomize each topic proportion to low and high proportion groups. The cutoffs are determined using the maximally selected rank statistics.
The AUC for each number of topics is shown in Figure 4(A). GpLSI achieves the highest area under the curve (AUC) at $K=6$ in both models. We also plot Kaplan Meier curves for each topic using the same dichotomized topic proportions. The result for GpLSI is illustrated in Figure 4(B). We observe that Topic 2, which is characterized by a high prevalence of granulocytes, and Topic 6, a mixture of CD4 T cells and blood vessels, are associated with lower cancer recurrence. Positive effect of granulocytes on cancer prognosis was also reported by Wu et al. (2022), who found out that a microenvironment with clustered granulocyte and tumor cells is associated with better patient outcomes. We observe the same association of granulocyte with lower risk in LDA Figure 16 of the Appendix.
4.2 Understanding Structure in Mouse Spleen Samples
Figure 5: (A) Visualization of estimated B cell microenvironment topics for $K=3,5,7,10$. (B) Comparison of clustering performance using Moran’s I and PAS score. We plot 1-PAS for better interpretation. (C) Estimated B cell microenvironment topic weights for $K=5$ using GpLSI.
We also apply our method to identify immune topics in mouse spleen. In this setting, each document is anchored to a B cell (Chen et al. 2020). A previous study has processed the original CODEX images from Goltsev et al. (2018) to obtain the frequencies of non-B cells in the 100 pixel neighborhood of each B cell (Chen et al. 2020). The final input for the topic models consists of a 35,271 B cell microenvironments by 24 cell types frequency matrix, along with the positional data of B cells.
We also evaluate the quality of clusters with two metrics, Moran’s I and the percentage of abnormal spots (PAS) (Shang & Zhou 2022). Moran’s I is a classical measure of spatial autocorrelation that assesses the degree to how values are clustered or dispersed across a spatial domain. PAS score measures the percentage of B cells for which more than 60% of its neighboring B cells have different topics. Higher Moran’s I and lower PAS score indicate more spatial smoothness of the estimated topics. From Figure 5(B), we conclude that GpLSI has the highest Moran I, and the lowest PAS scores, demonstrating improved spatial smoothness of the topics.
We observe that the B cell microenvironment topics identified with GpLSI align well with their biological context (Figure 5(C)). By referencing the manual annotations of B cells from the original study by Goltsev et al. (2018), we infer that Topic 1, Topic 2, Topic 3, and Topic 5 correspond to the red pulp, periarteriolar lymphoid sheath (PALS), B-follicles, and the marginal zone. This interpretation is further supported by high expression of CD4+T cells in Topic 2 (PALS) and high expression of marginal zone macrophages in Topic 5 (marginal zone).
4.3 Analysis of the “What’s Cooking” dataset
This dataset contains recipes from 20 different cuisines across Europe, Asia, and South America. Each recipe is a list of ingredients which allow us to convert to a count matrix with 13,597 recipes (documents) and 1,019 unique ingredients (words). Under the assumption that neighboring countries would have similar cuisine styles, we construct a graph of recipes based on the geographical proximity of the countries. Specifically, for each recipe, we select the five closest recipes from neighboring countries (including its own country) based on the $\ell_{1}$ distance of the ingredient count vectors and define them as neighboring nodes on the graph. Through this, we aim to identify general cooking styles that are prevalent across various countries worldwide.
We run GpLSI, pLSI, and LDA with $K=5,7,10,15,20$ topics. We illustrate the estimated topics of GpLSI for $K=7$ in Figure 6. The results for pLSI and LDA are provided in Section G of the Appendix. With this approach, Topic 1 is clearly a baking topic and Topic 6 is defined by strong spices and sauces common in Mexican or parts of Southeast Asian cuisines. We also observe a general topic for Asian cuisines (Topic 2) and another for Western countries (Topic 7). To evaluate the estimated topics, we compare each topic’s characteristics with the cuisine-by-topic proportion (Figure 6(C)). Indeed, the style of each topic defined by the anchor ingredients aligns with the cuisines that have a high proportion of that topic. For example, the baking topic (Topic 1) is prevalent in British, Irish, French, Russian, and South American cuisines.
In contrast, for pLSI, it is difficult to analyze the characteristics for each topic because Topics 1-4 have one or no identified anchor ingredients. Comparing the cuisine-by-topic proportions of GpLSI and LDA, we observe that GpLSI reveals many cuisines as mixtures of different cooking styles (Figure 6(B)). In contrast, for LDA, many cuisines such as Moroccan, Mexican, Korean, Chinese, Thai have their recipes predominantly classified to a single topic (Figure 18(B) of the Appendix). GpLSI provides estimates of topic mixture and topic weights that are more relevant to our goal of discovering global cooking styles.
5 Conclusion
We believe that our work offers valuable insights into structural topic models and opens up several avenues for further exploration. A promising direction is to incorporate structure to the topic matrix $A$ while jointly optimizing structural constraints on $W$. While our work focuses on low-$p$ regime, real world applications, such as genomics data with large $p$, could benefit from introducing sparsity to the word composition of each topic.
Appendix for “Graph Topic Modeling for Documents with Spatial or Covariate Dependencies”
Yeo Jin Jung and Claire Donnat
Department of Statistics, The University of Chicago
Appendix A Optimizing graph regularization parameter $\rho$
In this section, we propose a novel graph cross-validation method which effectively finds the optimal graph regularization parameter by partitioning nodes into folds based on a natural hierarchy derived from a minimum spanning tree. The procedure is summarized in the following algorithm.